OpenStudy (anonymous):

if i decrease the radius of earth to half. What is the effect on its gravitational force?

OpenStudy (anonymous):

According to the Newton's law of gravitation: \[F = \frac{ GMm}{ R^{2} }\]where G = gravitational constant. M = mass of Earth m = mass of a body on earth's surface R = distance between A and B from their centers. If the radius of the earth is halved, the gravitational force, \[F = \frac{ Gm_{1}m_{2} }{ (\frac{ 1 }{ 2 }R)^{2} } = \frac{ Gm_{1}m_{2} }{ \frac{ R^{2} }{ 4 } }\]Now, \[F = 4\frac{ GMm }{ R^{2} }\] Therefore, gravitational force will increase by four times.

OpenStudy (anonymous):

thanx bro

OpenStudy (anonymous):

you are welcomed...

OpenStudy (anonymous):

is it really R^2 in denominator

OpenStudy (anonymous):

yup! Force is inversely proportional to square of the distances from their centers.

OpenStudy (anonymous):

sorry, i have the formula potential diffrence in my mind. V=GMm/R

OpenStudy (anonymous):

its okay...